If I live 8 miles away from a store and it took me 30 minutes to run there and back in 20 minutes, what is the average speed? How do I find the average speed?

Answer by Alan3354(69443)

If I live 8 miles away from a store and it took me 30 minutes to run there and back in 20 minutes, what is the average speed? How do I find the average speed?
=============
That's 16 miles in 50 minutes.
d = r*t
r = d/t = 16/50
r = 0.32 miles/minute